# Magnon bands and transverse transport in a proposed two-dimensional $Cu_{2}F_{5}$ ferrimagnet

### v1

@@ 5 Conclusion
We have studied a 2D magnetic lattice which describes a layer of the $Cu_{2}F_{5}$ crystal [79, 80, 81] using the spin wave approach. The magnetic order is ferrimagnetic, with spin-up sites ($S=1$) pointing in the off-plane direction opposite to spin-down ($s=1/2$). The magnon bands are not generate, as expected for ferrimagnets and in contrast to AFM collinear order, where an effective time-reversal symmetry makes the bands degenerate. A non-null Berry curvature is generated by the Dzyaloshinskii-Moriya interaction between next-near neighbors. That enables the system to show transverse transport effects. An interesting remark is that, for high single-ion anisotropy ($A>J_{i}$), the directions of net spin (in response to a magnetic field gradient) and heat (in response to a temperature gradient) currents are sensitive to temperature, which can be technologically interesting regarding the control of magnon flow.

### v2

@@ 5 Conclusion
We have studied a 2D magnetic lattice that describes a layer of the $Cu_{2}F_{5}$ crystal [79, 80, 81] using the spin wave approach. The magnetic order is ferrimagnetic, with spin-up sites ($S=1$) pointing in the off-plane direction opposite to spin-down ($s=1/2$). The magnon bands are not degenerate, as expected for ferrimagnets, and in contrast to AFM collinear order in the Néel state, where an effective time-reversal symmetry makes the bands degenerate. A non-null Berry curvature is generated by the Dzyaloshinskii-Moriya interaction between next-near neighbors. That enables the system to show transverse transport effects. An interesting remark is that, for high single-ion anisotropy ($A\ \geq J_{2}$), the directions of net spin (in response to a magnetic field gradient) and heat (in response to a temperature gradient) currents are sensitive to temperature,resulting in non-monotonic curves for spin Hall coefficient and thermal Hall coefficient. That can be technologically interesting regarding the control of magnon flow. The non-monotonic curves seem to be an intrinsic feature of the system regardless of spin value, happening even when both spins are equal, provided that a single-ion anisotropy is present (and effective) and the magnon bands do not cross.